OpenStudy (anonymous):

Which of the following statements best describes the graph of 3x – 2y = 4?

OpenStudy (anonymous):

It is a straight line joining the points (6, 2), (5, 1), and (7, 3). It is a straight line joining the points (0, –2), (2, 1), and (–2, –5). It is a curve joining the points (3, –2), (2, 3), and (4, 1). It is a curve joining the points (0, 2), (–2, –1), and (4, 1).

OpenStudy (anonymous):

please help will fan and medal

OpenStudy (anonymous):

@Luigi0210 @zepdrix @amistre64 @SolomonZelman @undeadknight26

OpenStudy (anonymous):

could someone also help me with this one: For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 Answer choices: -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (triciaal):

from the info given standard form of eqn Ax + By = C 3x-2y = 4 it should be a line variable to the first power only when x = 0, y = -2 when y = 0 x = 4/3 y = -3/2(x) + 2

OpenStudy (triciaal):

correction y = 3/2(x) -2

OpenStudy (triciaal):

answer choice b (0,-2) (2,1) etc

OpenStudy (triciaal):

-x -2y =- 4 add x + 4 to each side -2y + 4 = x choice d

OpenStudy (anonymous):

thank you @triciaal

OpenStudy (triciaal):

welcome

OpenStudy (anonymous):

@tricaal what was the answer? D?